# MM-LLMs: Recent Advances in MultiModal Large Language Models

### 25 Jan 2024 v2

@@ Abstract
In the past year, MultiModal Large Language Models (MM-LLMs) have undergone substantial advancements, augmenting off-the-shelf LLMs to support MM inputs or outputs via cost-effective training strategies. The resulting models not only preserve the inherent reasoning and decision-making capabilities of LLMs but also empower a diverse range of MM tasks. In this paper, we provide a comprehensive survey aimed at facilitating further research of MM-LLMs. Specifically, we first outline general design formulations for model architecture and training pipeline. Subsequently, we provide brief introductions of existing MM-LLMs, each characterized by its specific formulations. Additionally, we review the performance of MM-LLMs on mainstream benchmarks and summarize key training recipes to enhance the potency of MM-LLMs. Lastly, we explore promising directions for MM-LLMs while concurrently maintaining a real-time tracking website111https://mm-llms.github.io for the latest developments in the field. We hope that this survey contributes to the ongoing advancement of the MM-LLMs domain.
@@ 1 Introduction
MM-LLMs harness LLMs as the cognitive powerhouse to empower various MM tasks. LLMs contribute desirable properties like robust language generation, zero-shot transfer capabilities, and In-Context Learning (ICL). Concurrently, foundation models in other modalities provide high-quality representations. Considering foundation models from different modalities are individually pre-trained, the core challenge facing MM-LLMs is how to effectively connect the LLM with models in other modalities to enable collaborative inference. The predominant focus within this field has been on refining alignment between modalities and aligning with human intent via a MM Pre-Training (PT) + MM Instruction-Tuning (IT) pipeline.
@@
With the debut of GPT-4(Vision) OpenAI (2023) and Gemini Team et al. (2023), showcasing impressive MM understanding and generation capabilities, a research fervor on MM-LLMs has been sparked. Initial research primarily focuses on MM content comprehension and text generation like (Open)Flamingo Alayrac et al. (2022); Awadalla et al. (2023), BLIP-2 Li et al. (2023c), Kosmos-1 Huang et al. (2023c), LLaVA/LLaVA-1.5 Liu et al. (2023e, d), MiniGPT-4 Zhu et al. (2023a), MultiModal-GPT Gong et al. (2023), VideoChat Li et al. (2023d), Video-LLaMA Zhang et al. (2023e), IDEFICS IDEFICS (2023), Fuyu-8B Bavishi et al. (2023), and Qwen-Audio Chu et al. (2023b). In pursuit of MM-LLMs capable of both MM input and output Aiello et al. (2023), some studies additionally explore the generation of specific modalities, such as Kosmos-2 Peng et al. (2023) and MiniGPT-5 Zheng et al. (2023b) introducing image generation, and SpeechGPT Zhang et al. (2023a) introducing speech generation. Recent research endeavors have focused on mimicking human-like any-to-any modality conversion, shedding light on the path to artificial general intelligence. Some efforts aim to amalgamate LLMs with external tools to reach an approaching ‘any-to-any’ MM comprehension and generation, such as Visual-ChatGPT Wu et al. (2023a), ViperGPT Surís et al. (2023), MM-REACT Yang et al. (2023), HuggingGPT Shen et al. (2023), and AudioGPT Huang et al. (2023b). Conversely, to mitigate propagated errors in the cascade system, initiatives like NExT-GPT Wu et al. (2023d) and CoDi-2 Tang et al. (2023b) have developed end-to-end MM-LLMs of arbitrary modalities. The timeline of MM-LLMs is depicted in Figure 1.
@@
In this paper, we present a comprehensive survey aimed at facilitating further research of MM-LLMs. To provide readers with a holistic understanding of MM-LLMs, we initially delineate general design formulations from model architecture (Section 2) and training pipeline (Section 3). We break down the general model architecture into five components: Modality Encoder (Section 2.1), Input Projector (Section 2.2), LLM Backbone (Section 2.3), Output Projector (Section 2.4), and Modality Generator (Section 2.5). The training pipeline elucidates how to enhance a pre-trained text-only LLM to support MM input or output, primarily consisting of two stages: MM PT (Section 3.1) and MM IT (Section 3.2). In this section, we also provide a summary of mainstream datasets for MM PT and MM IT. Next, we engage in discussions of State-of-the-Art (SOTA) MM-LLMs, each characterized by specific formulations, and summarize their development trends in Section 4. In Section 5, we comprehensively review the performance of major MM-LLMs on mainstream benchmarks and distill key training recipes to enhance the efficacy of MM-LLMs. In Section 6, we offer promising directions for MM-LLMs research. Moreover, we have established a website (https://mm-llms.github.io) to track the latest progress of MM-LLMs and facilitate crowd-sourcing updates. Finally, we summarize the entire paper in Section 7 and discuss related surveys on MM-LLMs in Appendix A. We aspire for our survey to aid researchers in gaining a deeper understanding of this field and to inspire the design of more effective MM-LLMs.
@@ 2.1 Modality Encoder
Various pre-trained encoder options exist for handling different modalities, where can be image, video, audio, 3D, or etc. Next, we will offer a concise introduction organized by modality.
@@
For images, there are generally four optional encoders: NFNet-F6 Brock et al. (2021), ViT Dosovitskiy et al. (2020), CLIP ViT Radford et al. (2021), and Eva-CLIP ViT Fang et al. (2023). NFNet-F6 is a normalizer-free ResNet He et al. (2016), showcasing an adaptive gradient clipping technique that allows training on extensively augmented datasets while achieving SOTA levels of image recognition. ViT applies the Transformer Vaswani et al. (2017) to images by first dividing the image into patches. It then undergoes linear projection to flatten the patches, followed by encoding via multiple Transformer blocks. CLIP ViT builds connections between text and images, comprising a ViT and a text encoder. Utilizing a vast amount of text-image pairs, it optimizes ViT by contrastive learning, treating paired text and images as positive samples and others as negative ones. Its Eva version stabilizes the training and optimization process of the massive CLIP, offering new directions in expanding and accelerating the expensive training of MM base models. For videos, they can be uniformly sampled to frames, undergoing the same pre-processing as images.
@@
is typically encoded by C-Former Chen et al. (2023b), HuBERT Hsu et al. (2021), BEATs Chen et al. (2023f), and Whisper Radford et al. (2023). C-Former employs the CIF alignment mechanism Dong and Xu (2020); Zhang et al. (2022a) for sequence transduction and a Transformer to extract audio features. HuBERT is a self-supervised speech representation learning framework based on BERT Kenton and Toutanova (2019), achieved by the masked prediction of discrete hidden units. BEATs is an iterative audio pre-training framework designed to learn Bidirectional Encoder representations from Audio Transformers.
@@
is typically encoded by ULIP-2 Salesforce (2022); Xu et al. (2023a, b) with a PointBERT Yu et al. (2022) backbone.
Moreover, to handle numerous heterogeneous modal encoders, some MM-LLMs, particularly any-to-any ones, use ImageBind Girdhar et al. (2023), a unified encoder covering six modalities, including image, video, text, audio, heat map, etc.
@@ 2.2 Input Projector
The Input Projector can be achieved directly by a Linear Projector or Multi-Layer Perceptron (MLP), i.e., several linear projectors interleaved with non-linear activation functions. There are also more complex implementations like Cross-attention, Q-Former Li et al. (2023c), or P-Former Jian et al. (2023). Cross-attention uses a set of trainable vectors as queries and the encoded features as keys to compress the feature sequence to a fixed length. The compressed representation is then fed directly into the LLM Bai et al. (2023b) or further used for X-text cross-attention fusion Alayrac et al. (2022). Q-Former extracts relevant features from , and the selected features are then used as prompts . Meanwhile, P-Former generates ‘reference prompts’, imposing an alignment constraint on the prompts produced by Q-Former. However, both Q- and P-Former require a separate PT process for initialization.
@@ 2.3 LLM Backbone
where the aligned representations of other modalities can be considered as soft Prompt-tuning for the LLM Backbone. Moreover, some research works have introduced Parameter-Efficient Fine-Tuning (PEFT) methods, such as Prefix-tuning Li and Liang (2021), Adapter Houlsby et al. (2019), and LoRA Hu et al. (2021). In these cases, the number of additional trainable parameters is exceptionally minimal, even less than 0.1% of the total LLM parameter count. We provide an introduction to mainstream PEFT methods in Appendix B.
The commonly used LLMs in MM-LLMs incude Flan-T5 Chung et al. (2022), ChatGLM Zeng et al. (2022a), UL2 Tay et al. (2022), Qwen Bai et al. (2023a), Chinchilla Hoffmann et al. (2022), OPT Zhang et al. (2022b), PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023a), LLaMA-2 Touvron et al. (2023b), and Vicuna Chiang et al. (2023). We provide a brief introduction to these models in Appendix C.
@@ 2.4 Output Projector
The optimization only relies on captioning texts, without utilizing any audio or visual resources , where and is the textual condition encoder in . The Output Projector is implemented by a Tiny Transformer or MLP.
@@ 2.5 Modality Generator
The Modality Generator is tasked with producing outputs in distinct modalities. Commonly, existing works use off-the-shelf Latent Diffusion Models (LDMs) Zhao et al. (2022), i.e., Stable Diffusion Rombach et al. (2022) for image synthesis, Zeroscope Cerspense (2023) for video synthesis, and AudioLDM-2 Liu et al. (2023b, c) for audio synthesis. The features mapped by the Output Projector serve as conditional inputs in the denoising process to generate MM content. During training, the ground truth content is first transformed into a latent feature by the pre-trained VAE Kingma and Welling (2013). Then, noise is added to to obtain the noisy latent feature . A pre-trained Unet Ronneberger et al. (2015) is used to compute the conditional LDM loss as follows:
optimize parameters and by minimizing .

### 20 Feb 2024 v4

@@ Abstract
In the past year, MultiModal Large Language Models (MM-LLMs) have undergone substantial advancements, augmenting off-the-shelf LLMs to support MM inputs or outputs via cost-effective training strategies. The resulting models not only preserve the inherent reasoning and decision-making capabilities of LLMs but also empower a diverse range of MM tasks. In this paper, we provide a comprehensive survey aimed at facilitating further research of MM-LLMs. Initially, we outline general design formulations for model architecture and training pipeline. Subsequently, we introduce a taxonomy encompassing MM-LLMs, each characterized by its specific formulations. Furthermore, we review the performance of selected MM-LLMs on mainstream benchmarks and summarize key training recipes to enhance the potency of MM-LLMs. Finally, we explore promising directions for MM-LLMs while concurrently maintaining a real-time tracking website111https://mm-llms.github.io for the latest developments in the field. We hope that this survey contributes to the ongoing advancement of the MM-LLMs domain.
@@ 1 Introduction
MM-LLMs harness LLMs as the cognitive powerhouse to empower various MM tasks. LLMs contribute desirable properties like robust language generation, zero-shot transfer capabilities, and In-Context Learning (ICL). Concurrently, foundation models in other modalities provide high-quality representations. Considering foundation models from different modalities are individually pre-trained, the core challenge facing MM-LLMs is how to effectively connect LLMs with models in other modalities to enable collaborative inference. The predominant focus within this field has been on refining alignment between modalities and aligning with human intent via a MM Pre-Training (PT) + MM Instruction-Tuning (IT) pipeline.
@@
With the debut of GPT-4(Vision) OpenAI (2023) and Gemini Team et al. (2023), showcasing impressive MM understanding and generation capabilities, a research fervor on MM-LLMs has been sparked. Initial research primarily focuses on MM content comprehension and text generation, encompassing tasks such as image-text understanding, exemplified by projects like BLIP-2 Li et al. (2023e), LLaVA Liu et al. (2023e), MiniGPT-4 Zhu et al. (2023a), and OpenFlamingo Awadalla et al. (2023); video-text understanding, as demonstrated by initiatives such as VideoChat Li et al. (2023f), Video-ChatGPT Maaz et al. (2023), and LLaMA-VID Li et al. (2023j); and audio-text understanding, as seen in projects like Qwen-Audio Chu et al. (2023b). Later, the capabilities of MM-LLMs have been expanded to support specific modality outputs. This includes tasks with image-text output, such as GILL Koh et al. (2023a), Kosmos-2 Peng et al. (2023), Emu Sun et al. (2024), and MiniGPT-5 Zheng et al. (2023b); as well as speech/audio-text output, exemplified by projects like SpeechGPT Zhang et al. (2023a) and AudioPaLM Rubenstein et al. (2023). Recent research endeavors have focused on mimicking human-like any-to-any modality conversion, shedding light on the path to artificial general intelligence. Some efforts aim to amalgamate LLMs with external tools to reach an approaching any-to-any MM comprehension and generation, such as Visual-ChatGPT Wu et al. (2023a), HuggingGPT Shen et al. (2023), and AudioGPT Huang et al. (2023b). Conversely, to mitigate propagated errors in the cascade system, initiatives like NExT-GPT Wu et al. (2023d), CoDi-2 Tang et al. (2023c), and ModaVerse Wang et al. (2024c) have developed end-to-end MM-LLMs of arbitrary modalities. The timeline of MM-LLMs is depicted in Figure 1.
@@
In this paper, we present a comprehensive survey aimed at facilitating further research of MM-LLMs. To provide readers with a holistic understanding of MM-LLMs, we initially delineate general design formulations from model architecture (Section 2) and training pipeline (Section 3). We break down the general model architecture into five components: Modality Encoder (Section 2.1), Input Projector (Section 2.2), LLM Backbone (Section 2.3), Output Projector (Section 2.4), and Modality Generator (Section 2.5). The training pipeline elucidates how to enhance a pre-trained text-only LLM to support MM input or output, primarily consisting of two stages: MM PT (Section 3.1) and MM IT (Section 3.2). In that section, we also provide a summary of mainstream datasets for MM PT and MM IT. Next, we establish a taxonomy encompassing State-of-the-Art (SOTA) MM-LLMs, each characterized by specific formulations, and summarize their development trends in Section 4. In Section 5, we comprehensively review the performance of major MM-LLMs on mainstream benchmarks and distill key training recipes to enhance the efficacy of MM-LLMs. In Section 6, we offer promising directions for MM-LLMs research. Moreover, we have established a website (https://mm-llms.github.io) to track the latest progress of MM-LLMs and facilitate crowd-sourcing updates. Finally, we summarize the entire paper in Section 7 and discuss related surveys on MM-LLMs in Appendix A. We aspire for our survey to aid researchers in gaining a deeper understanding of this field and to inspire the design of more effective MM-LLMs.
@@ 2.1 Modality Encoder
Various pre-trained encoder options exist for handling different modalities, where can be image, video, audio, 3D, etc. Next, we will offer a concise introduction organized by modality.
@@
For images, there are various optional encoders: NFNet-F6 Brock et al. (2021), ViT Dosovitskiy et al. (2020), CLIP ViT Radford et al. (2021), Eva-CLIP ViT Fang et al. (2023), BEiT-3 Wang et al. (2023d), OpenCLIP Cherti et al. (2023), Grounding-DINO-T Zhang et al. (2022b) with Swin-T Liu et al. (2021b) backbone, DINOv2 Oquab et al. (2023), SAM-HQ Kirillov et al. (2023) with MAE He et al. (2022), RAM++ Zhang et al. (2023i) with Swin-B backbone, InternViT Chen et al. (2023j), and VCoder Jain et al. (2023). For videos, they can be uniformly sampled to frames, undergoing the same pre-processing as images.
@@
Moreover, to handle numerous heterogeneous modal encoders, some MM-LLMs, particularly any-to-any ones, use ImageBind Girdhar et al. (2023), a unified encoder covering six modalities, including image/video, text, audio, heat map, inertial measurement units, and depth. We provide a brief introduction to some mainstream modality encoders in Appendix B.
@@ 2.2 Input Projector
The Input Projector can be achieved directly by a Linear Projector or Multi-Layer Perceptron (MLP), i.e., several linear projectors interleaved with non-linear activation functions. There are also more complex implementations like Cross-attention, Q-Former Li et al. (2023e), P-Former Jian et al. (2023), and MQ-Former Lu et al. (2023a). Cross-attention (Perceiver Resampler) Alayrac et al. (2022) uses a set of trainable vectors as queries and the encoded features as keys to compress the feature sequence to a fixed length. The compressed representation is then fed directly into the LLM or further used for X-Text cross-attention fusion. Q-Former extracts relevant features from with learnable queries, and the selected features are then used as prompts . Meanwhile, P-Former generates "reference prompts", imposing an alignment constraint on the prompts produced by Q-Former. MQ-Former conducts a fine-grained alignment of multi-scale visual and textual signals. However, both Q-, P-, MQ-Former require an additional PT process for initialization.
@@ 2.3 LLM Backbone
where the aligned representations of other modalities can be considered as soft Prompt-tuning for the LLM. Moreover, some works have introduced Parameter-Efficient Fine-Tuning (PEFT) methods, such as Prefix-tuning Li and Liang (2021), LoRA Hu et al. (2021), and LayerNorm tuning Zhao et al. (2024). In these cases, the number of additional trainable parameters is exceptionally minimal, even less than 0.1% of the total LLM parameter count. We provide an introduction to mainstream PEFT methods in Appendix C.
The commonly used LLMs in MM-LLMs incude Flan-T5 Chung et al. (2022), ChatGLM Zeng et al. (2022a), UL2 Tay et al. (2022), Persimmon Elsen et al. (2023), Qwen Bai et al. (2023a), Chinchilla Hoffmann et al. (2022), OPT Zhang et al. (2022c), PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023a), LLaMA-2 Touvron et al. (2023b), and Vicuna Chiang et al. (2023). We provide a brief introduction to some representative LLMs in Appendix D.
@@ 2.4 Output Projector
The optimization only relies on captioning texts, without utilizing any audio or visual resources , where and is the textual condition encoder in . The Output Projector is implemented by a Tiny Transformer with a learnable decoder feature sequence or MLP.
@@ 2.5 Modality Generator
The Modality Generator is tasked with producing outputs in distinct modalities. Commonly, existing works use off-the-shelf Latent Diffusion Models (LDMs) Song et al. (2021); Bao et al. (2022); Zhao et al. (2022), i.e., Stable Diffusion Rombach et al. (2022) for image synthesis, Zeroscope Cerspense (2023) for video synthesis, and AudioLDM-2 Liu et al. (2023b, c) for audio synthesis. The features mapped by the Output Projector serve as conditional inputs in the denoising process to generate MM content. During training, the ground truth content is first transformed into a latent feature by the pre-trained VAE Kingma and Welling (2013). Then, noise is added to to obtain the noisy latent feature . A pre-trained Unet Ronneberger et al. (2015) is used to compute the conditional LDM loss as follows:
which optimizes parameters and by minimizing .
@@ 3 Training Pipeline
forked edges,
for tree=
grow’=0,
draw,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
rounded corners,
align=left,
minimum width=4em,
edge+=darkgray, line width=1pt,
s sep=3pt,
inner xsep=2pt,
inner ysep=3pt,
line width=0.8pt,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=8em,font=,,
where level=2text width=5em,font=,,
where level=3text width=8em,font=,,
where level=4text width=5em,font=,,
MM-LLMs, ver
Functional Division
Underst.
I+TT: BLIP-2 Li et al. (2023e), Kosmos-1 Huang et al. (2023c), PaLM-E Driess et al. (2023),
ViperGPT
Surís et al. (2023), LLaVA Liu et al. (2023e), MiniGPT-4 Zhu et al. (2023a), mPLUG-Owl Ye et al. (2023b),
Otter Li et al. (2023b), MultiModal-GPT Gong et al. (2023),
PandaGPT Su et al. (2023), PaLI-X(Chen et al.)
LLaVA-Med Li et al. (2023d), LLaVAR Zhang et al. (2023h), mPLUG-DocOwl() Ye et al. (2023a), DLP
Jian et al. (2023), ChatSpot Zhao et al. (2023b), OpenFlamingo Awadalla et al. (2023), Chinese-LLaVA
LinkSoul-AI. (2023), ASM Wang et al. (2023c), BLIVA hu2 (2023), IDEFICS IDEFICS (2023), Qwen-VL
Bai et al. (2023b), Kosmos-2.5 Lv et al. (2023), InternLM-XComposer Zhang et al. (2023f), JAM
(Aiello et al.), LLaVA-1.5 Liu et al. (2023d), MiniGPT-v2 Chen et al. (2023d), Fuyu-8B Bavishi et al. (2023),
CogVLMWang et al. (2023b), mPLUG-Owl2 Ye et al. (2023c), Monkey Li et al. (2023l), Volcano
Lee et al. (2023), DRESS Chen et al. (2023i), LION Chen et al. (2023c), DocPedia() Feng et al. (2023),
ShareGPT4VChen et al. (2023f), VIM Lu et al. (2023b), mPLUG-PaperOwl()Hu et al. (2023a), RLHF-V
Yu et al. (2023b), Silkie Li et al. (2023g), Lyrics
Lu et al. (2023a), VILA Lin et al. (2023), CogAgent
Hong et al. (2023), Osprey Yuan et al. (2023a), V* Wu and Xie (2023), MobileVLM Chu et al. (2023a),
TinyGPT-V (Yuan et al.), DocLLM() Wang et al. (2023a), LLaVA- Zhu et al. (2024c), Yi-VLTeam. (2023)
KAM-CoT(Mondal et al.), InternLM-XComposer2 Dong et al. (2024b), MoE-LLaVA Lin et al. (2024a),
LLaVA-MoLE Chen et al. (2024),
LLaVA-NeXT Liu et al. (2024b), VLGuard Zong et al. (2024),
MobileVLM V2 Chu et al. (2024), ViGoRYan et al. (2024), VisLingInstruct Zhu et al. (2024b)
V+TT: VideoChat Li et al. (2023f), Video-ChatGPT Maaz et al. (2023), Dolphins Ma et al. (2023)
A+TT: SALMONN Tang et al. (2023a), Qwen-Audio Chu et al. (2023b)
3D+TT: 3DMIT Li et al. (2024b)
ManyT: Flamingo Alayrac et al. (2022), MM-REACT Yang et al. (2023b), X-LLM Chen et al. (2023b)
InstructBLIP Dai et al. (2023), EmbodiedGPT Mu et al. (2023), Video-LLaMA Zhang et al. (2023e), Lynx
Zeng et al. (2023), AnyMALMoon et al. (2023), LanguageBind Zhu et al. (2024a), LLaMA-VID
Li et al. (2023j), X-InstructBLIP Panagopoulou et al. (2023), InternVL Chen et al. (2023j)
, leaf, text width=46em
Generation
[ I+TI+T: FROMAGe() Koh et al. (2023b), Visual ChatGPT Wu et al. (2023a), DetGPT()Pi et al. (2023)
GILLKoh et al. (2023a), Kosmos-2() Peng et al. (2023), Shikra() Chen et al. (2023e), GPT4RoI()
Zhang et al. (2023g), SEED Ge et al. (2023), LISA() Lai et al. (2023), VisCPMHu et al. (2023b),
CM3LeonYu et al. (2023a), LaVIT Jin et al. (2024), DreamLLM Dong et al. (2024a), MiniGPT-5
Zheng et al. (2023b), Kosmos-G Pan et al. (2023), GLaMM() Rasheed et al. (2023), LLaVA-Plus(+&)
Liu et al. (2023f),
PixelLM() Ren et al. (2023), VL-GPT Zhu et al. (2023b), CLOVA(+&)
Gao et al. (2023b), Emu-2 Sun et al. (2023a), MM-Interleaved Tian et al. (2024), DiffusionGPT
Qin et al. (2024), RPGYang et al. (2024),Vary-toy() Wei et al. (2024), CogCoM() Qi et al. (2024),
SPHINX-X() Gao et al. (2024)
A/S+TA/S+T: SpeechGPT Zhang et al. (2023a), AudioPaLM Rubenstein et al. (2023)
ManyI+T: Emu Sun et al. (2024), BuboGPT() Zhao et al. (2023d), GroundingGPT() Li et al. (2024c)
ManyMany: GPT-4 OpenAI (2023), HuggingGPT Shen et al. (2023), AudioGPT Huang et al. (2023b)
NExT-GPT Wu et al. (2023d), ControlLLM Liu et al. (2023i), TEAL Yang et al. (2023a), CoDi-2(Tang et al.)
Gemini Team et al. (2023), ModaVerse Wang et al. (2024c), MLLM-ToolWang et al. (2024a)
, leaf, text width=46em
Design Division
Tool-using
[Visual ChatGPTWu et al. (2023a), ViperGPTSurís et al. (2023), MM-REACTYang et al. (2023b),
HuggingGPT Shen et al. (2023),AudioGPT Huang et al. (2023b), ControlLLM Liu et al. (2023i), LLaVA-Plus
(Liu et al.), CogAgent Hong et al. (2023),CLOVA Gao et al. (2023b), -UMi Shen et al. (2024), MLLM-Tool
(Wang et al.), WebVoyager He et al. (2024), Mobile-Agent Wang et al. (2024b)
, leaf, text width=46em
End-to-end
[The remaining models are essentially all end-to-end trainable models.
, leaf, text width=46em